OpenStudy (anonymous):

Could someone please help me with this question: Ascorbic acid (vitamin C) in water C6H8O6, 176.1 g/mol molarity = 1.92 M d = 1.22 g/mL molality = ?

OpenStudy (rogue):

Do you know what the answer is? I don't think its right, but I got 3.61 messing around... I can't seem to get my units to cancel.

OpenStudy (anonymous):

For the absorbic acid problem the answer is 2.18.

OpenStudy (rogue):

Alright, I'm over complicating it, hold on, give me a few minutes, I'll get the answer.

OpenStudy (rogue):

Yes! I got 2.17, but thats close enough. Let me post the work.

OpenStudy (rogue):

Ascorbic Acid: C6H8O2 Molar mass: 176.1 g/mol Molarity: 1.92 M = 1.92 mol/L Density: 1.22 g/mL = 1220 g/L From the density, 1 liter of this solution would weigh 1220 g. There would be 1.92 moles of ascorbic acid in 1 liter of this solution. 1.92 mol/L * 176.1 g/mol = 338.112 g/L C6H8O2 1220 g solution - 338.112 g C6H8O2 = 881.888 g H2O (solvent) = 0.881888 kg H2O Molality = moles solute / kg solvent \[m = \frac {1.92 mol C_6H_8O_2}{0.881888 kg H_2 O} \approx 2.17 mol/kg\]

OpenStudy (rogue):

Does that make sense?

OpenStudy (anonymous):

Yes- this makes sense now- thank you a ton, I have an exam on this topic tomorrow morning!

OpenStudy (rogue):

Alright, I wish you the best of luck! Ace that exam! :D

OpenStudy (anonymous):

Thank you- will do! :)